bobs boat travels at 12 mph in still water. find the speed of the current if he go 4 miles upstream in the same time that it takes to go 8 miles downstream. write an equation that could be used to model this problem. let c represent the speed (in mph) of the current stream.
Answer by macston(5194)

.
12mph+c=speed downstream, 12mph-c=speed upstream
.
4miles/(12mph-c)=8miles/(12mph+c) Cross-multiply.
96 mi^2/hr-8c miles=48mi^2/hr+4c miles Subtract 48mi^2/hr from each side
48mi^2/hr-8c miles=4c miles Add 8c miles to each side.
48mi^2/hr=12c miles Divide each side by 12 miles.
4 mi/hr=c
ANSWER: The rate of the current was 4 miles per hour.